# Max–Min Fairness in Stacked Intelligent Metasurface-Aided Rate Splitting Networks Abdullah Quran, Shimaa Naser, Maryam Tariq, Omar Alhussein, and Sami Muhaidat

### v1

@@ V Conclusion
In this paper, we investigated the integration of RSMA with a SIM for downlink multiuser MISO systems, focusing on a max–min fairness design. By leveraging SIM’s programmable wave-domain beamforming and 1-layer RSMA strategy, we formulated a max–min rate optimization problem and addressed its non-convexity via an AO framework. Specifically, the power allocation was handled through SCA, while SIM beamforming was optimized using the RCG method. Numerical results demonstrated that the proposed SIM-RSMA framework achieves higher minimum rates compared to SIM-NOMA and SIM-SDMA. These results demonstrate the potential of combining RSMA and SIM to enhance user fairness and maximize the minimum user rate in multiuser MISO systems.

### v2

@@ V Conclusion
In this paper, we investigated the integration of RSMA with a SIM for downlink multiuser MISO systems under a max–min fairness design. By leveraging SIM’s programmable wave-domain beamforming and the RSMA transmission strategy, we formulated and efficiently solved a max–min rate optimization problem using an alternating optimization framework, where power allocation was handled through successive convex approximation and SIM beamforming was optimized via the Riemannian conjugate gradient method. Simulation results showed that the proposed SIM-RSMA framework achieves performance nearly identical to SIM-SDMA while outperforming SIM-NOMA in terms of minimum user rate and fairness. These results indicate that, in the fully analog regime, the benefits of rate-splitting under max–min fairness are constrained by the shared SIM architecture and its unit-modulus phase-shift limitation, which restricts the independent shaping of the common stream and causes the system to effectively behave as SDMA. Therefore, future research should investigate hybrid or digitally assisted SIM-RSMA architectures that combine wave-domain beamforming with limited digital precoding, enabling user-specific control and fully exploiting RSMA’s interference management capability under fairness-driven objectives.